OpenStudy (anonymous):

Determine algebraically whether the function is even, odd, or neither even nor odd. 9/x^2

OpenStudy (sidsiddhartha):

do u know when a function is even and when it will be odd?

OpenStudy (anonymous):

If I see the graph I might

OpenStudy (sidsiddhartha):

yeah it can be done graphiclly and also algebrically let a function is f(x) then it will be even when \[f(x)=f(-x)\] and odd when \[f(-x)=-f(x)\] so in ur problem \[f(x)=\frac{ 9 }{ x^2 }\] now try to evaluate f(-x)

OpenStudy (anonymous):

I neither even nor odd

OpenStudy (sidsiddhartha):

how ???????? u have \[f(x)=\frac{ 9 }{ x^2 }\] so \[f(-x)=\frac{ 9 }{ (-x)^2 }=\frac{ 9 }{x^2 }\] so \[f(x)=f(-x)\] now what do u think even or odd?

OpenStudy (anonymous):

The function is even I think

OpenStudy (sidsiddhartha):

yes its even